# Generating Synthetic Electronic Health Record (EHR) Data: A Review with Benchmarking

### v1

@@ Abstract
Objectives: To conduct a scoping review of existing approaches for synthetic Electronic Health Records (EHR) data generation, to benchmark major methods with proposed open-source software, and to offer recommendations for practitioners.
@@
Results: 42 studies are identified and classified into five categories. Seven open-source methods covering all categories are selected, trained on MIMIC-III, and evaluated on MIMIC-III or MIMIC-IV for transportability considerations. Among them, Generative Adversarial Network (GAN)-based methods demonstrate competitive performance in fidelity and utility on MIMIC-III; rule-based methods excel in privacy protection. Similar findings are observed on MIMIC-IV, except that GAN-based methods further outperform the baseline methods in preserving fidelity. A Python package, “SynthEHRella”, is provided to integrate various choices of approaches and evaluation metrics, enabling more streamlined exploration and evaluation of multiple methods, which can be accessed at https://github.com/chenxran/synthEHRella.
Discussion: Method choice is governed by the relative importance of the evaluation metrics in downstream use cases. We provide a decision tree to guide the choice among the benchmarked methods.
@@
Keywords: Confidentiality, Diffusion Models, Generative Adversarial Network, Generative AI, Phecode, Privacy, Scoping Review, Synthetic EHR, Transformer, Variational Auto-Encoder.
@@ 1 Introduction
Electronic Health Records (EHRs) store patients’ health information collected via encounters with health systems in a digital format, consisting of both structured and unstructured data. EHRs are often linked to biorepositories that enable molecular profiling of blood or tissue samples. Such rich repositories that integrate disease phenotypes, multi-omics markers, medical images, medications, and laboratory results have become a critical resource for research in biomedicine [1, 2] and public health [3, 4, 5]. For researchers in medical informatics, bioinformatics, statistics, computer science, and other quantitative fields, EHR-linked biobanks [6, 7, 8] are valuable for empirical evaluations of analytical methods using multi-modal data [9, 10, 2, 11]. However, few real-world EHR data that are publicly available exist due to patient consent requirements and patient privacy considerations. Even when large databases are publicly available for research, when it comes to evaluation of methods through simulation studies, one needs repeated replicates of realistic EHR data.
@@
Synthetic EHR data generation has emerged as a promising solution to unlock enormous research and educational potential of real-world healthcare data while safeguarding data confidentiality. First, high-quality realistic synthetic EHR data enable the development and evaluation of analytical tools with less privacy concerns [12, 13, 14, 15, 16]. Second, for minority subpopulations, where the risk of reidentification is higher, using privacy-preserving synthetic data is ethicaly more defensible  [17, 18, 19]. Third, synthetic EHR generation facilitates experimenting with the size of the real training and testing data over multiple replicates for accuracy and uncertainty assessment of methods [20, 21]. Finally, synthetic data provide an ideal environment for training healthcare professionals, data scientists and other researchers in EHR usage, clinical workflows, and data analysis without the risk of exposing real patient information [22, 23].
@@
With the rapid growth in the literature on synthetic EHR generation (Figure S1), numerous review and discussion papers have focused on either recent methodological developments or on the drawbacks of existing evaluation metrics [24, 25, 26, 27, 28]. However, there is a notable lack of review papers with comprehensive benchmarking of existing methodologies using benchmark datasets. The latest benchmarking paper [29], while making valuable contributions to evaluating Generative Adversarial Network (GAN)-based [30] synthetic EHR generation methods, has two major limitations. First, it evaluated exclusively the methods that are built around GAN, falling short of offering a more comprehensive comparison against alternative approaches in other categories we identified in Section 3.1. Second, the benchmarking was based on closed-source data, limiting the reproducibility and the feasibility for evaluating future methods by other researchers. Finally, with an explosion in latest generative artificial intelligence (GenAI) methods, including large language models (LLMs) and diffusion models, a more comprehensive and updated review, evaluation, and benchmarking is needed.
In this paper, we address this knowledge gap by conducting a scoping review of the extant literature on synthetic EHR generation methods and provide a comprehensive benchmarking of seven current approaches and two baseline approaches. Our study provides a holistic evaluation of synthetic EHR generation methods using multiple evaluation metrics, with open-source datasets to ensure reproducibility and accessibility for future research. We also provide an open-source Python package “SynthEHRella” that can be readily used on-premise for synthetic EHR methods evaluation.
Based on our literature review, we classify existing methods into five categories: Rule-based, GAN-based, Variational Autoencoder (VAE)-based [31], Transformer-based [32], and Diffusion-based methods [33, 34]. We include 1-2 representative methods from each category that have usable publicly-available codes and compare their performances on three critical aspects: fidelity, downstream utility, and privacy [29, 27, 35]. Specifically, fidelity assesses how closely the synthetic data resemble the real EHR data distributions, often in terms of summary features such as the number of visits per patient or the marginal distribution of International Classification of Disease (ICD) codes or PhecodeX [36], an extended version of the Phecode [37], which is built from human-crafted grouping of ICD codes [^1] We will use “phecode” or “phecodeX” interchangeably to refer to the codes in the PhecodeX system.. Downstream utility evaluates the gap between using synthetic data and the real data when performing downstream tasks. Privacy examines the exposure risk of sensitive patient information (e.g., their inclusions in the dataset, their clinical attributes) in synthetic EHR data. In addition to the quality of synthetic data, we also evaluate the computational time cost associated with each method.
Most prior studies focused on using the MIMIC-III dataset [38] for training the data generation engines and for evaluation. We expand our evaluation to include the newly released MIMIC-IV dataset [39]. The inclusion of the MIMIC-IV dataset enables us not only to evaluate the performance of the existing methods on a larger patient cohort with more recent data, but to examine how would the existing methods that have been trained and evaluated on MIMIC-III perform under a different but related population, namely to assess the transportability of the synthetic data generation methods under distributional shifts. Based on our evaluation results, we provide recommendations for practitioners and identify areas where further developments are required.
We summarize the key contributions of this paper in the following:
  1. 1.
    We present a scoping review and comprehensive, multi-pronged benchmarking of synthetic EHR generation methods on open-source datasets, considering fidelity, utility, privacy, and computational cost;
  2. 2.
    We incorporate the latest MIMIC-IV dataset into our benchmarking, in addition to MIMIC-III, to provide an evaluation of existing methods under potential distributional shift in the training and testing populations to assess transportability;
  3. 3.
    Based on our study, we present a decision tree to offer recommendations to practitioners. We identify key gaps in current approaches and indicate potential avenues for future research;
  4. 4.
    We develop and release the SyntEHRella package to facilitate and encourage future benchmarking and reproducibility efforts in synthetic EHR generation.
The rest of the paper is organized as follows. In Section 2, we introduce methods for literature review and inclusion criteria of the methods included in our benchmarking. We then discuss data source in Section 2.2, and introduce the evaluation procedure, including statement of the evaluation problem, settings, and metrics in Section 2.3. Numerical studies on the effect of varying input training dataset sizes or the generated synthetic dataset sizes are presented in Section 2.4. In Section 3, we present our findings. Specifically, we summarize our results for literature review in Section 3.1; we describe the summary statistics of the data used for benchmarking in Section 3.2; we then present our main findings on evaluation and numerical studies in Section 3.3 and Section 3.4, respectively. We introduce the open-source SynthEHRella package in Section 3.5. In Section 4, we interpret our findings, present a decision tree based on our current evaluation results to offer recommendations for practitioners, discuss challenges of existing methods, outline areas for future methods development, and limitations of this work. The paper concludes with remarks in Section 5.
@@ 2 Method
2.1 Literature Review
2.1.1 Search Procedure
In the spirit of a scoping review, we conducted a search on May 5th, 2024 to identify published and preprint articles in English that propose methods for generating synthetic EHR data. The search keywords include (a) synthetic EHR, (b) synthetic EHR generation, and (c) synthetic electronic health records. The search was performed across three academic databases: Google Scholar, PubMed, and Semantic Scholar. An additional search verification was conducted on August 30th, 2024 to identify updated literature.
All search results underwent a rigorous screening by the authorship team to exclude irrelevant or unqualified papers. First, studies that did not focus on pipeline development of synthetic EHR generation are removed. Second, papers without an available codebase for reproducing or generating synthetic EHR data were excluded from the consideration of benchmarking. In our final slate, we chose seven distinct methods from each of the five categories that are frequently present in the literature. We also added two baseline methods, one based on directly resampling from the real dataset not intended for privacy-sensitive use but for anchoring the values of the performance metrics, and another based on generating binary phecodes using independent Bernoulli distributions with the mean equal to the marginal phenotype prevalence of each phecode. Detailed descriptions of these methods and the baselines are provided in Section 3.1.
@@ 2.2 Benchmarking Datasets
In this work, the evaluation is carried out on two open-source large-scale Intensive Care Unit (ICU)-based and emergency department (ED)-based [^2] ED-based admissions appear only in the MIMIC-IV dataset. benchmark EHR datasets described as follows:
MIMIC-III [38]
@@
MIMIC-III is a large-scale EHR dataset comprising 58,976 unique ICU admissions at Beth Israel Deaconess Medical Center (BIDMC) in Boston, Massachusetts between 2001 and 2012. The dataset consists of mainly adult patients, as well as 7,874 neonates admitted to the Neonatal Intensive Care Unit (NICU). All patient data have been de-identified to ensure patient privacy. The dataset uses International Classification of Disease codes (ICD-9) to record diagnostic events. Many existing studies on synthetic EHR data have focused their evaluations primarily on the presence or count of codes for diagnoses, medications, and procedures within the MIMIC-III dataset.
MIMIC-IV [39]
@@
MIMIC-IV is a recently released, updated version of MIMIC series datasets that features a larger patient cohort, using both version 9 and 10 of ICD as coding system. MIMIC-IV includes 431,231 hospital admissions of patients that were admitted into either the emergency department or ICUs between 2008 and 2019 at BIDMC. Due to the overlap of data collection periods (2008-2012), a subset of patients in MIMIC-IV also appear in MIMIC-III, although it is difficult, if not impossible, to link profiles across datasets because of the de-identification procedures. As a recently-released dataset, few studies have evaluated their synthetic data methods on MIMIC-IV. To the best of our knowledge,  [40, 41, 42] are the only existing works that have assessed their methods on this dataset.
In this study, version 1.4 of MIMIC-III and version 2.2 of MIMIC-IV are used for the comparative evaluation and benchmarking. To highlight the difference from “synthetic data”, we use “real data” to refer to either the MIMIC-III or MIMIC-IV data which we will distinguish when necessary; we use “training data” to refer to the subset of data for training methods which can be real, synthetic data, or their combination. Note that the scenario where the synthetic data can be treated as training data only appear in the predictive utility evaluation described in Section 2.3.
2.3 Evaluation
2.3.1 Statement of Evaluation Task
@@
In this study, we focus on the generation of cross-sectional diagnostic events data, following the capabilities of most publicly-available codebases. ICU-based (and ED-based) EHR datasets, such as MIMIC-III and MIMIC-IV, have a longitudinal nature, where each individual patient may have multiple encounters. To this end, we convert all longitudinal records into a cross-sectional format by aggregating each patient’s records into a single representation. Formally, consider an EHR dataset with cross-sectional format $\mathcal{D}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$, where each entry $\mathbf{x}_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(K)})^{\top}$ is a $K$-dimensional binary vector representing the presence/absence of diseases coded by a specific coding system (e.g., ICD-9). Each dimension $x_{i}^{(k)}=1$ if the $k$-th disease is diagnosed at least once in all visits of $i$-th patient, $k=1,\cdots,K$, $i=1,\cdots,N$. The goal of synthetic EHR generation is to produce a synthetic dataset of size $M$, $\mathcal{D}_{syn}=\{\mathbf{x}_{i,syn}\}^{M}_{i=1}$, where each sample in $\mathcal{D}_{syn}$ can be viewed as samples drawn from the same distribution as those in the real dataset of size $N$ denoted as $\mathcal{D}_{real}=\{\mathbf{x}_{i,real}\}^{N}_{i=1}$, while ensuring privacy protection for patients in $\mathcal{D}_{real}$. We represent the datasets by two $N\times K$ and $M\times K$ binary matrices denoted as $\mathbf{X}_{real}$ and $\mathbf{X}_{syn}$, respectively. For this study, we fix $M=50,000$ except that in the numerical studies (see Section 2.4) we vary size of generated samples to investigate their impacts on the quality of synthetic data.
The entire MIMIC-III dataset is used to train generative models if no pre-trained models are available. For a fair evaluation, we adopt PhecodeX 1.0 [36] as the standard coding system for evaluation. Before evaluation, all synthetic EHR data are first mapped to the ICD-10 coding system, and then to PhecodeX. Specifically, we map ICD-9 to ICD-10 based on the icd-mappings package in Python [^3] https://pypi.org/project/icd-mappings/0.1.1/; we map Systematized Medical Nomenclature for Medicine–Clinical Terminology (SNOMED-CT) to ICD-10 via the SNOMED mapping tool [^4] https://prod-mapping.ihtsdotools.org/#/project/records; we then convert ICD-10 coding system to PhecodeX according to the Phecode Map X (Extended) [^5] https://phewascatalog.org/phecode_x. Based on MIMIC-III, we identify $K=2,254$ distinct phecodes. For visualization, we keep only phecodes with more than 50 patients in MIMIC-III, resulting in $K=1,773$ phecodes. For quantitative evaluation, on the other hand, all additional granularity (i.e., digits after the decimal point) in phecodes were omitted and merged into their corresponding parent categories, resulting in $K=595$ parent phecodes.
2.3.2 Performance Evaluation Metrics
We evaluate the quality of synthetic EHR generated by all the selected methods and baselines on both MIMIC-III and MIMIC-IV from three widely adopted perspectives: 1) fidelity for evaluating how closely the distribution of synthetic EHR data matches with real EHR data; 2) downstream utility for measuring the performance gap in downstream analyses compared to real data; 3) privacy exposure for assessing how well the synthetic EHR generation methods protect patient information. Additionally, we also evaluate the selected methods’ computational efficiency. The evaluation metrics are presented below.
@@ 2.3.2 Performance Evaluation Metrics
To evaluate whether the generative algorithms capture phenotype-level prevalence information, we first evaluate the dimension-wise distributional discrepancy between the real and the synthetic datasets. We compute the Maximum Mean Discrepancy (MMD) to evaluate the absolute difference in prevalence. Similarly, we use Root Mean Squared Percentage Error (RMSPE) and Mean Absolute Percentage Error (MAPE) for assessing the relative difference in prevalence. These metrics are computed as follows:
@@
$\texttt{RMSPE}=100\times\sqrt{\frac{1}{K}\sum_{i=1}^{K}\left(\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right)^{2}};$
$\texttt{MAPE}=\frac{100}{K}\sum_{i=1}^{K}\left|\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right|.$
All the RSMPE values are divided by a factor of 100 for ease of presentation.
In addition to quantitative evaluation, we also present visualizations to compare the distribution characteristics between the real and the synthetic EHR data. Specifically, we use a boxplot to show the dimension-wise prevalence using each method, and another boxplot to show the distribution of the number of unique phecodes per patient in the data generated by each method. A scatterplot is included to compare the dimension-wise prevalence between real and synthetic EHR data, with each dot representing a phecode.
We are also interested in evaluating whether generative models preserve pair-wise correlations between the phecodes. To this end, we calculate the Pearson correlation matrix following [29, 43] for both real and synthetic datasets and compute the correlation Frobenius distance (CFD):
@@
where $a^{real}_{kk^{\prime}}$ and $a^{syn}_{kk^{\prime}}$ represents the ($k,k^{\prime}$)-th element in correlation matrices $\hat{\Sigma}_{real}$ and $\hat{\Sigma}_{syn}$, respectively. In addition, we visualize element-wise difference of correlation between the real and the synthetic data via boxplots; 100,000 pairs of entries from the upper triangle of the two correlation matrices are sampled and shown in the boxplots.
Finally, we assess whether we can distinguish the synthetic and the real data used to train the synthetic data generation method, referred to as discriminative prediction. We use both the real and the synthetic data to train a logistic regression model for predicting whether a given sample in the combined data is real or synthetic. We validate the model via 5-fold cross-validation. We report the Area Under the Curve (denoted as AUC) and Accuracy (denoted as ACC) of the classifier, where lower values indicate higher resemblance of the synthetic data to the real data.
@@
A primary objective of generating synthetic EHR data is to replace the use of real data in downstream applications, thereby protecting patient privacy without compromising data utility. To this end, we evaluate the utility of synthetic data by examining their performance in both analytical and predictive tasks.
Firstly, analytical utility assesses the capability of the synthetic data in providing inferential results that approximate those produced based on the real data. To this end, we focus on the analytical task of estimating associations between binary phenotypes $x_{i}^{(k)}$ and $x_{i}^{(k^{\prime})}$ by fitting a logistic regression model with the $k$-th phenotype as the outcome, and the $k^{\prime}$-th phenotype as the explanatory variable. The point estimates and the 95% confidence intervals of regression coefficients (i.e., log-odds ratios) of the predictor phenotype are reported. A larger overlap in the 95% Confidence Intervals (CIs) between those obtained based on the real data and the synthetic data indicates higher analytical utility of the synthetic data generation method.
Secondly, predictive utility measures the capability of the synthetic data in training machine learning (ML) methods for downstream use. To evaluate existing methods’ predictive utility, a binary classification task is constructed, where the $k$-th phenotype $x^{(k)}_{i}$ is treated as the outcome with the remaining phenotypes in $\mathbf{x}_{i}$ as predictors. In this work, we list three scenarios commonly used in the literature [29, 44] that concern different combinations of the sources of data used for training and testing a ML method. All the testing are based on the real data for objective evaluations. Suppose we split the real dataset into training and testing subsets (4:1 ratio). All the three scenarios are visualized in Figure S2, with detailed descriptions as follows:
  1. 1.
  2. 2.
    “Train on Synthetic + Real, Test on Real” (TSRTR). This setting investigates if the synthetic data can be considered as a data augmentation of the real data in downstream applications which have been shown to be capable of boosting performance by preventing overfitting issues when only a limited amount of real data is available [45] We stack the training split in the real data and the synthetic data together to train the ML model and test it on the test set of the real data. In practice, this is only feasible when the investigator has access to the real data. Nonetheless, it is included here for showing whether synthetic data has deleterious or beneficial effect when added to a real data;
  3. 3.
    “Train on Real and Test on Real”(TRTR). This scenario does not use synthetic data. The ML model is trained on the training data split, and find the test error on the testing data split. This provides the baseline in predictive utility evaluation.
For each of the above three scenarios, we calculate the AUC and ACC of the classifiers. Because TRTR is a real data based approach, we use it as a baseline and report the differences: ${\rm AUC}_{TSTR}-{\rm AUC}_{TRTR}$ and ${\rm AUC}_{TSRTR}-{\rm AUC}_{TRTR}$, and similarly for ACC. Higher values indicate better utility of the synthetic data.
In this study, analytical utility evaluations are conducted under two tasks. In the first task, whether a patient gets any type of cancer (phecodeX starting with “CA”) is treated as the outcome, and Obesity (EM_236) is treated as the predictor; for the second task, Diabete (EM_202) is treated as the outcome, and Hypertension (CV_401) is treated as the predictor. See Table 2 for prevalences of any type of cancers, obesity, hypertension, and diabetes. For predictive utility evaluation, Hypertension (CV_401) is treated as the outcome, because it is the most prevalent phenotype in both MIMIC-III and MIMIC-IV cohorts; the remaining phenotypes are treated as predictors. The logistic regression model is treated as our ML models, with results of more advanced ML methods, including Random Forest [46], Gradient Boosting [47], XGBoost [48], LightGBM [49], and K-Nearest Neighbours (K-NN), reported in Supplementary Materials.
@@
First, we use membership inference risk (MIR) [50] to assess the risk that an attacker, who gains access to a patient’s complete medical record, could determine whether this patient is included in the training dataset. A subject’s membership to the training dataset may be sensitive because the patient may wish to keep this information private from potential attackers. For example, the training data may be based on all the HIV-positive patients treated in a regional facility. Latest work of membership inference attack using synthetic data proposed to conduct privacy attack via an attacker model that can detect local overfitting which occurs if the synthetic data density evaluated at the real data record has a high value; this typically occurs when the real data record is close to the bulk of synthetic data [51]. We compute the minimum Euclidean distance between each real medical record and the synthetic EHR dataset:
@@
where $\mathbf{x}_{i,real}$ is a record from the real dataset. We further provide an average characterization by reporting the mean and median of these minimum distances as the dataset-level risk of membership inference if all the subjects in the real data may be attacked:

### v2

# Generating Synthetic Electronic Health Record Data: a Methodological Scoping Review with Benchmarking on Phenotype Data and Open-Source Software

@@ Abstract
Objectives: To conduct a scoping review of existing approaches for synthetic Electronic Health Records (EHR) data generation, to benchmark major methods, and to provide an open-source software and offer recommendations for practitioners.
@@
Results: 48 studies are identified and classified into five categories. Seven open-source methods covering all categories are selected, trained on MIMIC-III, and evaluated on MIMIC-III or MIMIC-IV for transportability considerations. Among them, Generative Adversarial Network (GAN)-based methods demonstrate competitive performance in fidelity and utility on MIMIC-III; rule-based methods excel in privacy protection. Similar findings are observed on MIMIC-IV, except that GAN-based methods further outperform the baseline methods in preserving fidelity.
Discussion: Method choice is governed by the relative importance of the evaluation metrics in downstream use cases. We provide a decision tree to guide the choice among the benchmarked methods. An extensible Python package, “SynthEHRella”, is provided to facilitate streamlined evaluations.
@@ 1 INTRODUCTION
Electronic Health Records (EHRs) store patients’ health information collected via encounters with health systems in a digital format. Such rich repositories that integrate disease phenotypes, multi-omics markers, medical images, medications, and laboratory results have become a critical resource for research in biomedicine[fodeh2016mining, yadav2018mining] and public health[birkhead2015uses, friedman2013electronic, kruse2018use]. For researchers in quantitative biomedical areas, EHR-linked biobanks[sudlow2015uk, all2019all, zhou2022global] are valuable for empirical evaluations of analytical methods using multi-modal data[tarczy2013survey, linder2021role, yadav2018mining, beesley2020emerging]. However, few real-world EHR data that are publicly available exist due to patient consent requirements and patient privacy considerations.
@@
Synthetic EHR data generation offers a promising solution to unlock enormous research and educational potential of real-world healthcare data while safeguarding confidentiality. First, high-quality synthetic EHR data enable the development and evaluation of analytical tools with less privacy concerns[rankin2020reliability, murcia2024automating, shi2022generating, perets2023subpopulation, muller2022synthesising]. Second, it is ethically more defensible[zhang2020ensuring, biswal2021eva, bing2022conditional] for minority subpopulation with higher re-identification risk. Third, synthetic EHR generation facilitates experimenting with the size of the real training and testing data over multiple replicates for accuracy and uncertainty assessment of methods[tall2020generating, chen2022simulation]. Finally, synthetic data provide ideal risk-free environments for training healthcare professionals, data scientists and other researchers in EHR usage, clinical workflows, and data analysis[yale2020generation, laderas2017teaching].
@@
Despite the growing literature on synthetic EHR generation (Figure S1) along with numerous review papers[mendelevitch2021fidelity, hernandez2022synthetic, ghosheh2022review, budu2024evaluation, achterberg2024evaluation], few works have conducted comprehensive benchmarking of existing methodologies. The latest benchmarking paper[yan2022multifaceted], while making valuable contributions to benchmarking synthetic EHR generation methods, has two major limitations. First, it evaluated exclusively the methods built around Generative Adversarial Network (GAN), overlooking other categories of approaches we identified in Section 3.1. Second, it relied on closed-source data, limiting the reproducibility and future evaluation by other researchers. With an explosion in latest generative artificial intelligence (GenAI) methods, including large language models (LLMs) and diffusion models, a more comprehensive and updated review, evaluation, and benchmarking is needed.
This paper makes four key contributions to address this knowledge gap. First, we conduct a methodological scoping review to summarize existing works for methodological development in synthetic EHR generation. Second, we propose a benchmarking on fidelity, utility, and privacy evaluation of diverse categories of methods that were not included in previous reviews and benchmarking studies[yan2022multifaceted, budu2024evaluation, goncalves2020generation]. Third, we use open-source EHR phenotype datasets, MIMIC-III and the recently released MIMIC-IV datasets, as our benchmarks and release SynthEHRella, an open-source, extensible, on-premises benchmarking toolkit not available in prior studies. Finally, we evaluate the transportability capability of existing methods by training them on MIMIC-III and evaluate on MIMIC-IV. Two comparison tables regarding the methods and evaluation metrics included in our study and the latest benchmarking work[yan2022multifaceted], Table S5 and Table S6, are presented in the Supplementary Materials. The comprehensive scoping review in this paper and the software will be relevant for EHR researchers and the medical informatics community moving forward.
@@ 2 METHOD
2.1 Scoping review
In the spirit of a scoping review, we conducted a search on May 5th, 2024 to identify studies that propose synthetic EHR generation methods. The full search strings are provided in Section S1 of the Supplementary Materials. Searches were performed on Google Scholar, PubMed, and Semantic Scholar. For PubMed, we used the built-in search engine, while for the others we employed the Publish or Perish software[harzing2007pop] to retrieve papers. Non-English papers, patents, and non-research outputs were excluded via manual and automated filtering. Two additional verification searches were conducted on August 30th, 2024 and March 30th, 2025, separately to identify any updated literature before August 30th, 2024.
All search results underwent a rigorous screening by the authorship team to exclude papers unrelated to synthetic EHR pipeline development. For the inclusion-eligible literature, the following information was extracted: algorithm name, model type, year of publication, support for longitudinal and conditional generation, evaluation metrics used, codebase accessibility, whether evaluation was conducted on a MIMIC dataset, and, if so, which data modalities were included. We refer the readers to the PRISMA-ScR checklist[tricco2018prisma] in the Supplementary Materials for details.
The results of the scoping review are concluded in Section 3.1 and Table 1. In terms of the benchmarking evaluation, we selected seven distinct methods based on the specific guiding principles. We refer the readers to Section 2.3 and 3.2 for details.
@@ 2.2 Benchmarking datasets
We introduce our benchmarking datasets, with detailed description of the datasets, coding system, and descriptive analysis in Section S1.
MIMIC-III[johnson2016mimic]
@@
MIMIC-III is a large-scale EHR dataset comprising 58,976 unique ICU admissions at Beth Israel Deaconess Medical Center (BIDMC) in Boston, Massachusetts between 2001 and 2012. The dataset consists of mainly adult patients, and 7,874 neonates admitted to the Neonatal Intensive Care Unit (NICU). All patient data have been de-identified for patient privacy. The dataset uses International Classification of Disease codes (ICD-9) to record diagnostic events.
MIMIC-IV[johnson2023mimic]
@@
MIMIC-IV is a recently released, updated version of MIMIC series datasets that features a larger patient cohort, using both version 9 and 10 of ICD as coding system. MIMIC-IV includes 431,231 hospital admissions of patients that were admitted into either the emergency departments (ED) or ICUs between 2008 and 2019 at BIDMC. Due to the overlap of data collection periods (2008-2012), a subset of patients in MIMIC-IV also appear in MIMIC-III, although it is difficult to link profiles across datasets because of the de-identification procedures.
2.3 Benchmarking evaluation
Statement of Evaluation Task
@@ 2.3 Benchmarking evaluation
ICU-based (and ED-based) EHR datasets, such as MIMIC-III and MIMIC-IV, have a longitudinal nature, where each individual patient may have multiple encounters. To this end, we convert all longitudinal records into a cross-sectional format by aggregating each records into a representation. Formally, consider an EHR dataset with cross-sectional format $\mathcal{D}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$, where each entry $\mathbf{x}_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(K)})^{\top}$ is a $K$-dimensional binary vector representing the presence/absence of diseases coded by a specific coding system (e.g., ICD-9). Each dimension $x_{i}^{(k)}=1$ if the $k$-th disease is diagnosed at least once in all visits of $i$-th patient, $k=1,\cdots,K$, $i=1,\cdots,N$. We represent the datasets by two $N\times K$ and $M\times K$ binary matrices denoted as $\mathbf{X}_{real}$ and $\mathbf{X}_{syn}$, respectively.
Evaluation Metrics
We introduce key evaluation metrics. Full descriptions, including the evaluation of computational cost and numerical studies, are presented in Section S1.
@@
We first evaluate the dimension-wise distributional discrepancy between the real and the synthetic datasets. We compute the Maximum Mean Discrepancy (MMD) to evaluate the absolute difference in prevalence. We use Root Mean Squared Percentage Error (RMSPE) and Mean Absolute Percentage Error (MAPE) for assessing the relative difference in prevalence. They are computed as follows:
@@
$\texttt{RMSPE}=100\times\sqrt{\frac{1}{K}\sum_{i=1}^{K}\left(\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right)^{2}},\leavevmode\nobreak\ \texttt{MAPE}=\frac{100}{K}\sum_{i=1}^{K}\left|\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right|.$
All the RMSPE values are divided by a factor of 100 for ease of presentation.
To evaluate whether generative models preserve pair-wise correlations between the phecodes, we calculate the Pearson correlation matrix following[yan2022multifaceted, yuan2023ehrdiff] for both real and synthetic datasets and compute the correlation Frobenius distance (CFD):
@@
where $a^{real}_{kk^{\prime}}$ and $a^{syn}_{kk^{\prime}}$ represent the ($k,k^{\prime}$)-th element in correlation matrices $\hat{\Sigma}_{real}$ and $\hat{\Sigma}_{syn}$, respectively.
Finally, we assess whether synthetic data can be distinguished from the real data used for training, a task referred to as discriminative prediction. We train a logistic regression model on both real and synthetic data to predict whether a given sample is real or synthetic. We do not fine-tune pre-trained checkpoints for discriminators, which was applied in existing work[zhang2022keeping], due to fair comparison consideration. See Section S1 for detailed explanations. We validate the model via 5-fold cross-validation. We report the Area Under the Curve (AUC) and Accuracy (ACC), where lower values indicate greater similarity between synthetic and real data.
@@
We evaluate the downstream utility of the synthetic data from two aspects. Firstly, analytical utility assesses how well synthetic data approximates inferential results derived from real data. Specifically, we estimate associations between binary outcome and explanatory phenotypes, denoted by $x_{i}^{(k)}$ and $x_{i}^{(k^{\prime})}$, by fitting a logistic regression model. The point estimates and the 95% confidence intervals of regression coefficients (i.e., log-odds ratios) of the predictor phenotype are reported. Greater overlap in 95% Confidence Intervals (CIs) from real and synthetic data indicates higher analytical utility.
Secondly, predictive utility measures how effectively synthetic data trains predictive machine learning (ML) models for downstream tasks. To evaluate existing methods’ predictive utility, we construct a binary classification task, treating the $k$-th phenotype $x^{(k)}_{i}$ as the outcome, and the remaining phenotypes in $\mathbf{x}_{i}$ as predictors. In this work, we list three scenarios commonly used in the literature[yan2022multifaceted, li2023generating] that concern different combinations of the sources of data used for training and testing a ML method: a) Train on Synthetic, Test on Real (TSTR); b) Train on Synthetic + Real, Test on Real (TSRTR); c) Train on Real and Test on Real (TRTR). All the testings are based on a same subset of real data for fair comparison. All the three scenarios are visualized in Figure S2, with detailed description in Section S1.
For each of the above three scenarios, we calculate the AUC and ACC of the classifiers. Since TRTR is real-data-based, we use it as a baseline and report the differences: ${\rm AUC}_{TSTR}-{\rm AUC}_{TRTR}$ and ${\rm AUC}_{TSRTR}-{\rm AUC}_{TRTR}$, and similarly for ACC. Higher values indicate better utility of the synthetic data.
@@
First, we evaluate membership inference risk (MIR)[shokri2017membership] which evaluates the likelihood that an attacker with access to a patient’s complete medical record, could determine whether this patient was included in the training dataset. We compute the minimum Euclidean distance between each real medical record and the synthetic EHR dataset:
@@
where $\mathbf{x}_{i,real}$ is a record from the real dataset. To quantify dataset-level membership inference risk, we report the mean and median of these minimum distances: